5) Find the Partial Decomposition of each rational expression. 5x-2 b) (a2-4)(x+1) b) 23+2x2+x
Answer #1

Answer & = s + 5x-2 (2-3) (ara)(22+1) D (An+B) + 2+1 Simplify right hand side 5x-2 = (Ar+B) 6-2) (7172) + C 662 +) 12+2)+ (x-yeti A+C+ D =0 B+ 2 C-2D=0 - 4 A + C + D - 5 -4Brac-2D = -2 __ on solving these equations - A = -1, B = 2 c = z D = 3 Therefo(2+2) - (A+B) 2? + (2 A+B+C)* PA. A+B= 0 0 2A+B+C = 1–0 TA=2 - from equation -2 B=-A = B=-2 And from equation ④ 2622-1+0=1 2i hope you got your answer but if you need any further clarification in any step then let me know through comment section